# Cryptomite: A versatile and user-friendly library of randomness extractors

### v1

@@ 4.2 Cryptomite  for Randomness Extraction in (Q)RNG
In random number generation (RNG), randomness extraction is used to process the raw output from an entropy source (sometimes called the noise source) into the final output that is $\epsilon$-perfectly random – a process analogous to conditioning in the NIST standards [57]. In most protocols for quantum RNG (QRNG), this is done using (strong) seeded randomness extractors. To showcase Cryptomite for RNG, we replicate and extend the randomness extraction step of the semi-device independent QRNG based on heterodyne detection [54], again giving code examples.
In [54], the randomness extractor error is chosen to be $\epsilon=10^{-10}$ and the length of the extractor input is $n_{1}=6.5\times 10^{9}$ bits [^11] Again, since the extractor input length is bigger than $2^{29}$, the extractor implementation will use the bigNTT, implying a throughput of 34×3-4$\times$3 - 4 × less than that shown in Section 2.3.. The experimental results and calculations certify a min-entropy rate of $k_{1}/n_{1}=0.08943$ and, therefore, a total min-entropy of $k_{1}=\lfloor 0.08943n_{1}\rfloor=581295000$ for the extractor input. The randomness extraction step is performed using the $\mathsf{Toeplitz}$ extractor, which gives an output length of $m=581294933$ bits and requires a seed length of $n_{2}=n_{1}+m-1=7081294934$ bits. This can be performed in a few lines of code using Cryptomite, as shown in Figure 7.
Extension – As with the QKD example, the RNG case can be improved using other options given in our extractor library, reducing the seed size and allowing the option to relax the requirement on the seed min-entropy rate. Using our $\mathsf{Circulant}$ extractor instead of the $\mathsf{Toeplitz}$ extractor saves $581294932$ seed bits. There is a small cost to this, as one needs to adjust the input length $n_{1}=6.5\times 10^{9}$ to $n_{1}^{\prime}$ such that $n_{1}^{\prime}+1$ is prime and adjust the extractor inputs min-entropy (using either Theorem 8 or Lemma 1), leading to a new output length of $581294865$ bits (i.e. $68$ less output bits). [^12] Note that the output of $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ is in principle 1 bit longer than $\mathsf{Circulant}$. In this example, the closest prime to $n_{1}$ is $62$ bits closer than the closest prime with primitive root 2 – so this in principle output length gain is eliminated in practice, and one gets a longer output using $\mathsf{Circulant}$. Moreover, using a seeded extractor for random number generation demands a (near-)perfectly random seed as the resource, which leads to a circularity. As in the QKD example, using Cryptomite we introduce a min-entropy rate parameter $r$ that allows for the extractor seed to be weakly random only, whilst still outputting near-perfect randomness. The code for this extension can be found in Figure 8.
@@
%**** main.tex Line 900 ****def improved_randomness_extraction(raw_randomness: list, seed_bits: list,
@@
# To use Circulant, the seed length needs to be a prime,
# we find the one closest to n_1 + 1 (see Def. 9):
n_adjusted = cryptomite.utils.closest_prime(n_1+1)
@@
# Compute the extractors output length (product sources
@@ 5 Conclusion and Future Work
  • We are in the process of implementing other randomness extractors, such as an efficient (in quasi-linear complexity) version of the Raz’ [4] extractor and the ones of Hayashi et al. [8]. These may be released in subsequent versions of Cryptomite.
  • The $\mathsf{Trevisan}$ extractor boasts an asymptotically small seed length. However, in practice it has the drawback that its computation time is large and, for small input lengths, the seed length is often longer than the input. Some interesting future work would be making a GPU implementation of $\mathsf{Trevisan}$, so that the one-bit extraction step is parallelised and fast in practice. One could alternatively find a combinatorial weak design that allows for a short seed length for small input lengths, whilst retaining minimal overlap between the chunks (see Section 3.6).
@@
  • To prove that some of our extractors are quantum-proof in the product sources model, we employ the proof techniques of [36] and references therein which rely on obtaining a bound on the collision entropy of the extractor with a uniform seed. This bound is well understood for extractors based on two-universal hashing families. It would be interesting to find a generalised theorem that allows for the security of any two-source extractor in the quantum-proof product sources model (through a bound on the collision entropy of a general extractor or otherwise).
@@ 6 Acknowledgements
We thank Kevin Milner and Kieran Wilkinson for valuable feedback and comments, Sean Burton for reviewing and improving the Cryptomite code, Dan Neville for a useful code review of our Trevisan extractor, Sherilyn Wright for designing the repository logo and Lluis Masanes for useful discussions. We also acknowledge Ela Lee and Matty Hoban for testing the first version of Cryptomite.
@@ Appendix A Appendices
A.1 The Number Theoretic Transform and the convolution theorem
The $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors in Cryptomite can be expressed as convolutions, allowing an efficient implementation by using the Fast Fourier Transform (FFT) or Number Theoretic Transform (NTT). However, the FFT uses floating point arithmetic, which may cause loss of precision that is unacceptable, for example, in cryptographic applications. Instead, we use the NTT with modular arithmetic to implement the same ring operations without precision issues.
Definition 10 (Number Theoretic Transform (NTT)).
The number theoretic transform on a vector $v$ of length $n$ is a discrete Fourier transform computed using the ring of integers modulo $p$ ($\mathbb{Z}/p\mathbb{Z}$) instead of the ring of complex numbers ($\mathbb{C}$) with primitive root $\omega$ (instead of $e^{-i\frac{2\pi}{n}}$) such that $\omega^{n}=1$. The inverse NTT, $\mathsf{NTT}^{-1}$, is defined such that $\mathsf{NTT}^{-1}(\mathsf{NTT}(v))=v$ and
@@ A.1 The Number Theoretic Transform and the convolution theorem
$\mathsf{NTT}(v*w)_{i}$ $=\sum_{k}^{n-1}\left(\sum_{j=0}^{n-1}{v_{j}\cdot v_{k-j}}\right)\omega^{ik}$ (51)
$=\sum_{j=0}^{n-1}{v_{j}\omega^{ij}\sum_{k=0}^{n-1}{v_{k-j}\omega^{i(k-j)}}}$ (52)
$=\left(\sum_{j=0}^{n-1}{v_{j}\omega^{ij}}\right)\left(\sum_{t=0}^{n-1}{v_{t}\omega^{it}}\right)\qquad\text{Let $t=k-j$}$ (53)
$=\mathsf{NTT}(v)_{i}\cdot\mathsf{NTT}(v)_{i}$ (54)
@@
The convolution theorem allows for the implementation of convolution in $O(n\log n)$ time using the NTT (or FFT), rather than the naive method requiring $O(n^{2})$. This efficiency improvement makes randomness extraction feasible for sizes $n>10^{6}$, as shown in Section 2.3 “Performance”.
A.2 Proofs for the $\mathsf{Circulant}$ extractor
@@
In order to prove that our $\mathsf{Circulant}$ construction is an extractor with the parameters specified in Section 3, we must first prove that the family of functions, constructed from Circulant matrices, is a two-universal family of hash functions. Once this is shown, we can exploit the well-known Leftover Hash Lemma [49] and use Corollary 5.6.1 from [17] to prove the construction is directly quantum-proof. One of the difficulties in proving that the hash functions from Circulant matrices are two-universal arises from a set of ‘problematic’ inputs (concretely, any pair of inputs that differ in every position), which we avoid by padding the extractor’s input. This can be achieved with a single constant bit padding, in turn minimising the required amount of additional seed. We start by stating some useful Lemmas and Theorems.
@@
For any prime $n$ and $x^{\prime}\in\{0,1\}^{n}$, such that $x^{\prime}\neq\{0\}^{n},\{1\}^{n}$, the Circulant matrix generated by $x^{\prime}$, $\mathrm{circ}(d^{\prime})$, is non-singular.
@@
The proof follows directly from considering a special case of Proposition 23 in [65], where elements of $x^{\prime}$ are bits, and noting that $x^{\prime}\neq\{0\}^{n},\{1\}^{n}$, by definition. ∎
Lemma 2.
Any square non-singular matrix is a bijection.
Proof.
This property follows directly from the invertibility of non-singular matrices. ∎
Lemma 3.
Let a random variable $Y\in\{0,1\}^{n}$ be such that $p_{Y}(y)=2^{-n}$ for all $y$. Then, for any non-singular $n\times n$ matrix $D$, $p_{DY}(y)=2^{-n}$ for all $y$.
Proof.
For each $y\in\{0,1\}^{n}$, there is only one value $Y=y^{*}$ s.t. $Dy^{*}=y$, and this value is unique from the fact that $D$ is a bijection. This means that, if
$p_{Y}(y)=2^{-n}\hskip 28.45274pt\forall y,$ (55)
then
$p_{DY}(y)=p_{Y}(y^{*})=2^{-n}\hskip 28.45274pt\forall y$ (56)
Lemma 4.
Given a random variable $Y\in\{0,1\}^{n}$ such that $p_{Y}(y)=2^{-n}$ for all $y\in\{0,1\}^{n}$, the first $m$ bits of $Y$, denoted $Y_{0:m-1}$ satisfy
$p_{Y_{0:m-1}}(u)=2^{-m}$ (57)
for all $u\in\{0,1\}^{m}$.
Proof.
We use that $p_{Y_{0:m-1}}(u)$ are marginals:
$p_{Y_{0:m-1}}(u)$ $=\sum\limits_{v\in{\{0,1\}^{n-m}}}p_{Y_{0:m-1},Y_{m:n-1}}(u,v)$ (58)
$=\sum\limits_{v\in{\{0,1\}^{n-m}}}p_{Y}([u,v])=2^{-n}2^{n-m}=2^{-m}$ (59)
Theorem 16.
The function in Definition 9 is a two-universal family of hash functions.
Proof.
The extractor function ($\mathsf{Circulant}$) in Definition 9 is a two-universal family of hash functions if, for all $x\neq\tilde{x}\in\{0,1\}^{n-1}$ and $Y\in\{0,1\}^{n}$ such that $p_{Y}(y)=2^{-n}$ for all $y$, we have
$P\left(\mathsf{Circulant}(x,Y)=\mathsf{Circulant}(\tilde{x},Y)\right)\leq 2^{-m}.$ (60)
Let $x^{\prime}=[x,0]$, $\tilde{x}^{\prime}=[\tilde{x},0]$ and define $d^{\prime}=x^{\prime}\oplus\tilde{x}^{\prime}$, where $\oplus$ denotes element wise addition modulo 2. Then, we can re-write the probability as
$P\left(\mathsf{Circulant}(x,Y)=\mathsf{Circulant}(\tilde{x},Y)\right)$ $=P((\mathrm{circ}(d^{\prime})Y)_{0:m-1}=\{0\}^{m})=p_{(\mathrm{circ}(d^{\prime})Y)_{0:m-1}}(\{0\}^{m}).$ (61)
Next, $x,\tilde{x}$ differ by at-least one bit (by definition), so $d^{\prime}\neq\{0\}^{n}$ and $x^{\prime},\tilde{x}^{\prime}$ share the same last bit (from the padding), so $d^{\prime}\neq\{1\}^{n}$. Then, by Theorem 15, $\mathrm{circ}(d^{\prime})$ is non-singular, and since $Y$ is uniformly distributed, we can use Lemma 3 to bound the probability $p_{\mathrm{circ}(d^{\prime})Y}(y)$ for all $y\in\{0,1\}^{n}$. Specifically, this leads to
$p_{\mathrm{circ}(d^{\prime})Y}(\{0\}^{n})=2^{-n}$ (62)
which, in combination with Lemma 4, gives
$p_{(\mathrm{circ}(d^{\prime})Y)_{0:m-1}}(\{0\}^{m})=2^{-m}$ (63)
and completes our proof. ∎
@@
A.3 $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ (and $\mathsf{Circulant}$) implementation
Concretely, our implementation of the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor follows [16] with $\{A_{i}\}_{i}$ chosen to be right cyclic shift matrices, which means loosing one output bit compared the optimal choice for the extractor. We also require that $n$, the input length, is prime with primitive root 2 and $X\neq\{0\}^{n},\{1\}^{n}$. The set of matrices $A_{1},A_{2}\ldots A_{n}$ are defined as the $n\times n$ matrices
$A_{1}=\begin{bmatrix}1&0&0&\ldots&\ldots&0\\ 0&1&0&\ldots&\ldots&0\\ 0&0&1&\ldots&\ldots&0\\ \vdots&&\ddots&&\vdots\\ 0&0&0&\ldots&1&0\\ 0&0&0&\ldots&0&1\\ \end{bmatrix},\quad A_{2}=\begin{bmatrix}0&1&0&\ldots&\ldots&0\\ 0&0&1&\ldots&\ldots&0\\ 0&0&0&\ldots&\ldots&0\\ \vdots&&\ddots&&\vdots\\ 0&0&0&\ldots&0&1\\ 1&0&0&\ldots&0&0\\ \end{bmatrix},\ldots,$ (76)
and satisfy the constraint that the sum of any subset $B\subset\{A_{i}\}_{i}$ has rank at-least $n-1$, as proved in [66]. In this form, the extractor can be re-written as the matrix-vector multiplication
$\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}(x,y)=\left(\begin{bmatrix}x_{0}&x_{1}&x_{2}&\ldots&\ldots&x_{n-1}\\ x_{n-1}&x_{0}&x_{1}&\ldots&\ldots&x_{n-2}\\ x_{n-2}&x_{n-1}&x_{0}&\ldots&\ldots&x_{n-3}\\ \vdots&&\ddots&\ddots&&\vdots\\ x_{2}&x_{3}&x_{4}&\ldots&x_{0}&x_{1}\\ x_{1}&x_{2}&x_{3}&\ldots&x_{n-1}&x_{0}\\ \end{bmatrix}\begin{bmatrix}y_{0}\\ y_{1}\\ \vdots\\ \vdots\\ y_{n-1}\end{bmatrix}\right)_{0:m},$ (88)
where $x=x_{0},x_{1},\ldots x_{n-1}$, $y=y_{0},y_{1},\ldots y_{n-1}$ and the subscript $1:m$ denotes the first $m$ elements (bits) of the matrix-vector multiplication.
Although this construction is not optimal in the sense that it loses one output bit compared to the optimal construction, it allows for the overall function to be rewritten as a circular convolution ([16], Appendix D.1, Definition 10). This means that we can write
@@
where subscript $i\in\{1,\ldots m\}$ denotes the $i$th output, $\mathsf{R}(x)=x_{0},x_{n-1},x_{n-2},\ldots,x_{1}$, and indices are mod $n$. The output consists of the first $m$ elements of this convolution. The extractor algorithm is given by the pseudo-code in Algorithm 1.
@@
[…]
2:     $L\leftarrow 2^{\text{floor}(\log_{2}(2n-2)+1)}$triangleright Calculate NTT size

### v3

@@ 5.2 Cryptomite  for randomness extraction in (Q)RNG
In random number generation (RNG), randomness extraction is used to process the raw output from an entropy source (sometimes called the noise source) into the final output that is $\epsilon$-perfectly random – a process analogous to conditioning in the NIST standards [68]. In most protocols for quantum RNG (QRNG), this is done using (strong) seeded randomness extractors. To showcase Cryptomite for RNG, we replicate and extend the randomness extraction step of the semi-device-independent QRNG based on heterodyne detection [65], again giving code examples.
In [65], the randomness extractor error is chosen to be $\epsilon=10^{-10}$ and the length of the extractor input is $n_{1}=6.5\times 10^{9}$ bits [^12] Again, since the extractor input length is bigger than $2^{29}$, the extractor implementation will use the bigNTT, implying a throughput of $3-4\times$ less than that shown in Section 2.3.. The experimental results and calculations certify a min-entropy rate of $k_{1}/n_{1}=0.08943$ and, therefore, a total min-entropy of $k_{1}=\lfloor 0.08943n_{1}\rfloor=581295000$ for the extractor input. The randomness extraction step is performed using the $\mathsf{Toeplitz}$ extractor, which gives an output length of $m=581294933$ bits and requires a seed length of $n_{2}=n_{1}+m-1=7081294932$ bits. This can be performed in a few lines of code using Cryptomite, as shown in Figure 7.
Extension – As in the QKD example, the RNG case can be improved using other options given in our extractor library, reducing the seed length and allowing the option to relax the requirement on the seed min-entropy rate. Using our $\mathsf{Circulant}$ extractor instead of the $\mathsf{Toeplitz}$ extractor saves $581294863$ seed bits. To implement this, the input length $n_{1}=6.5\times 10^{9}$ must be adjusted to $n_{1}^{\prime}$ such that $n_{1}^{\prime}+1$ is a prime number with primitive root 2. Based on this adjustment to $n_{1}$, the min-entropy must also be recalculated (using either Theorem 8 or Lemma 1). We find that the closest prime with primitive root 2 is $6.5\times 10^{9}+69$, so we set $n_{1}^{\prime}=6.5\times 10^{9}+68$ and pad the weak input with $68$ fixed bits. Moreover, using a seeded extractor for random number generation demands a (near-)perfectly random seed as the resource, which leads to a circularity. As in the QKD example, using Cryptomite we introduce a min-entropy rate parameter $r$ that allows the extractor seed to be weakly random only, whilst still outputting near-perfect randomness. The code for this extension can be found in Figure 8.
@@
def improved_randomness_extraction(raw_randomness: list, seed_bits: list,
@@
# To use Circulant, the seed length needs to be a prime
# with primitive root 2, we find the one closest
# to n_1 + 1 (see Def. 9):
n_adjusted = cryptomite.utils.closest_na_set(n_1+1)
@@
# Compute the extractors output length (product source
@@ 6 Conclusion and future work
  • We are in the process of implementing other randomness extractors, such as an efficient (in quasi-linear computation time) version of the Raz’ extractor [6] ([73]) and the ones of Hayashi et al. [7]. These may be released in subsequent versions of Cryptomite.
  • The $\mathsf{Trevisan}$ extractor boasts an asymptotically small seed length. However, in practice it has the drawback that its computation time is large and, for small input lengths, the seed length is often longer than the input [^13] We remark that the seed length can be made small by reducing the output length, i.e. at the cost of the protocol’s efficiency. See the exact statements of the seed length in Section 4.4. Some interesting future work would be making a GPU implementation of $\mathsf{Trevisan}$, so that the one-bit extraction step is parallelised and fast in practice. One could alternatively find a combinatorial weak design that allows for a short seed length for small input lengths, whilst retaining minimal overlap between the chunks (see Section 4.4).
@@
  • To prove that some of our extractors are quantum-proof in the product source model, we employ the proof techniques of [46] and references therein which rely on obtaining a bound on the collision entropy of the extractor with a uniform seed. This bound is well understood for extractors based on two-universal hashing families. It would be interesting to find a generalised theorem that allows for the security of any two-source extractor in the quantum-proof product source model (through a bound on the collision entropy of a general extractor or otherwise).
Since the completion of this work, we have been made aware of [74], in which the authors’ prove that the family of $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractors is quantum-proof in the Markov model with substantially better parameters than those obtained by the generic reduction of [13].
@@ 7 Acknowledgements
We thank Kevin Milner and Kieran Wilkinson for valuable feedback and comments, Sean Burton for reviewing and improving the Cryptomite code, Dan Neville for a useful code review of our Trevisan extractor, Sherilyn Wright for designing the repository logo, Ela Lee and Matty Hoban for testing the first version of Cryptomite, and Lluis Masanes for useful discussions. We also thank Martin Sandfuchs for pointing out that the generalisation to the Markov model (Theorem 4 in [13]) is unnecessary for extractors that are already quantum-proof in the product source model, enabling us to derive Corollary 1, and Christoph Pacher for pointing out a mistake in the original proof of Theorem 15.
@@ Appendix A Appendices
A.1 Comparison of seeded extractors
Seeded extractors are used for numerous applications. Trevisan’s extractor has the asymptotically shortest seed length of the seeded extractors implemented in this work. However, this is not always the case in practise. In this subsection, we provide a figure demonstrating that the asymptotic performance is not always preserved for short input lengths. We consider only those that have quantum-proof in the seeded case with near minimal entropy loss (i.e., not $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$).
Figure 9: A plot displaying the seed length of our extractor algorithms as a function of the input length ($n_{1}$) for each seeded extractor of Cryptomite. The seed length is calculated by fixing the input min-entropy to $k_{1}=\frac{n_{1}}{2}$ and calculating the maximum output length according to Figure 3.
We remark that the seed length of the $\mathsf{Trevisan}$ extractor depends on the output length $m$ and the extractor error $\epsilon$. For small $m$, the seed will be shorter – a situation that naturally occurs in the case of $k_{1}\ll n_{1}$ and may be relevant in certain experimental situations. Alternatively, the seed length can be further reduced by using a different configuration for the weak design, at the cost of additional entropy loss. For a more detailed explanation, see Section 4.4.
A.2 The number-theoretic transform and the convolution theorem
The $\mathsf{Circulant}$, $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ and $\mathsf{Toeplitz}$ extractors in Cryptomite can be expressed as convolutions, allowing an efficient implementation by using the fast Fourier transform (FFT) or number-theoretic Transform (NTT). However, the FFT uses floating point arithmetic, which may cause loss of precision that is unacceptable, for example, in cryptographic applications. Instead, we use the NTT with modular arithmetic to implement the same ring operations without precision issues.
Definition 10 (Number-theoretic transform (NTT)).
The number-theoretic transform on a vector $v$ of length $n$ is a discrete Fourier transform computed using the ring of integers modulo $p$ ($\mathbb{Z}/p\mathbb{Z}$) instead of the ring of complex numbers ($\mathbb{C}$) with primitive root $\omega$ (instead of $e^{-i\frac{2\pi}{n}}$) such that $\omega^{n}=1$. The inverse NTT, $\mathsf{NTT}^{-1}$, is defined such that $\mathsf{NTT}^{-1}(\mathsf{NTT}(v))=v$ and
@@ A.2 The number-theoretic transform and the convolution theorem
$\mathsf{NTT}(v*w)_{i}$ $=\sum_{k}^{n-1}\left(\sum_{j=0}^{n-1}{v_{j}\cdot w_{k-j}}\right)\omega^{ik}$ (47)
$=\sum_{j=0}^{n-1}{v_{j}\omega^{ij}\sum_{k=0}^{n-1}{w_{k-j}\omega^{i(k-j)}}}$ (48)
$=\left(\sum_{j=0}^{n-1}{v_{j}\omega^{ij}}\right)\left(\sum_{t=0}^{n-1}{w_{t}\omega^{it}}\right)\qquad\text{Let $t=k-j$}$ (49)
$=\mathsf{NTT}(v)_{i}\cdot\mathsf{NTT}(w)_{i}$ (50)
@@
The convolution theorem allows for the implementation of convolution in $O(n\log n)$ time using the NTT (or FFT), rather than the naive method requiring $O(n^{2})$. This efficiency improvement makes randomness extraction feasible for input lengths $n>10^{6}$, as shown in Section 2.3 “Performance”.
A.3 Proofs for the $\mathsf{Circulant}$ extractor
@@
To prove that our $\mathsf{Circulant}$ construction is an extractor with the parameters in Section 4, we first show that the family of functions derived from circulant matrices forms a two-universal hash family. With this established, we apply the Leftover Hashing Lemma [60] and Corollary 5.6.1 of [19] to conclude that the construction is quantum-proof. The main difficulty in proving the two-universality of the $\mathsf{Circulant}$ extractor comes from certain ‘problematic’ input pairs (those differing in every position). To resolve this, we pad the extractor input with a single constant bit, eliminating problematic cases while minimising additional seed length. We now state some useful Lemmas and Theorems.
Lemma 2 ([75]).
Let $n$ be a prime with 2 as a primitive root modulo $n$. Let $d^{\prime}\in\{0,1\}^{n}\setminus\{\{0\}^{n},\{1\}^{n}\}$ be any binary vector of length $n$ that is not all zeros or all ones. Let $\mathrm{circ}(d^{\prime})\in\{0,1\}^{n\times n}$ be the circulant matrix generated by $d^{\prime}$, where the $i$th row, for $i\in\{0,\ldots,n-1\}$, is $d^{\prime}$ right-shifted by $i$ positions. Then the rank of $\mathrm{circ}(d^{\prime})$ over $\{0,1\}$ satisfies $\mathrm{rank}(\mathrm{circ}(d^{\prime}))\geq n-1$.
Lemma 3.
Let $C=\mathrm{circ}(d^{\prime})\in\{0,1\}^{n\times n}$ be the circulant matrix generated by $d^{\prime}\in\{0,1\}^{n}$, where $n$ is a prime such that 2 is a primitive root modulo $n$. Suppose $d^{\prime}\notin\{\{0\}^{n},\{1\}^{n}\}$. Then any submatrix of $C$ formed by selecting $m\leq n-1$ rows has full row rank over $\{0,1\}$.
Proof.
By Lemma 2, $C=\mathrm{circ}(d^{\prime})$ has at-least rank n1n-1 over $\{0,1\}$. When $C$ has full rank, the Lemma is trivially true, so suppose that $C\in\{0,1\}^{n\times n}$ has $\mathrm{rank}(C)=n-1$. In this case, the space spanned by the rows of $C$ has dimension n1n-1, implying there exists a unique (up to scaling) linear dependence among the $n$ rows. We can write this, representing the $i$th row as $r_{i}\in\{0,1\}^{n}$, as
$\bigoplus_{i=0}^{n-1}\alpha_{i}r_{i}=\{0\}^{n},\quad\text{for some }\boldsymbol{\alpha}=(\alpha_{0},\ldots\alpha_{n-1})\in\{0,1\}^{n},\ \boldsymbol{\alpha}\neq\{0\}^{n},$ (51)
where the elementwise row addition is done modulo 2.
For any matrix with rows $\{r_{i}\}_{i}$ and a linear dependency vector $\boldsymbol{\alpha}$, any permutation of the columns preserves the linear dependency. In the case of a circulant matrix, shifting each column to the right by $j$ positions (cyclically) is equivalent to shifting each row element, for all rows, to the right by $j$ positions, resulting in the row that was originally $j$ rows below. Specifically, for any $j\in\{0,\ldots,n-1\}$, the $i$th row after cyclically permuting the columns by $j$ positions becomes $r_{(i+j)\bmod n}$. This means that we have
$\bigoplus_{i=0}^{n-1}\alpha_{i}r_{i}=\bigoplus_{i=0}^{n-1}\alpha_{i}r_{(i+j)\bmod n}=\{0\}^{n},$ (52)
for any $j\in\{0,\ldots,n-1\}$. By relabelling the coefficients accordingly, this expression is equivalent to
$\bigoplus_{i=0}^{n-1}\alpha_{i}r_{i}=\bigoplus_{i=0}^{n-1}\alpha_{(i-j)\bmod n}r_{i}=\{0\}^{n}.$ (53)
The second term corresponds to a linear dependency given by $j$ left cyclic shifts of $\boldsymbol{\alpha}$, implying that any cyclic shift of a dependency $\boldsymbol{\alpha}$ is also a dependency.
Putting this together: (i) since $\mathrm{rank}(C)=n-1$, the null space of $C$ is one-dimensional, so any nonzero vector $\boldsymbol{\alpha}$ in the null space is unique up to scalar multiplication; and (ii) since $C$ is circulant, any cyclic shift of a dependency vector is also a dependency. The only non-all-zero vector in $\{0,1\}^{n}$ that is invariant under all cyclic shifts and unique up to scalar multiplication is the all-ones vector, i.e., $\boldsymbol{\alpha}=(1,1,\dots,1)$. Thus, the only linear dependency among the rows of $C$ is $\sum_{i=0}^{n-1}r_{i}=\{0\}^{n}$.
This dependency involves all $n$ rows, so any submatrix formed by selecting $m\leq n-1$ rows must have full row rank (i.e., rank $m$) over $\{0,1\}$. This completes the proof. ∎
Lemma 4.
Let $Y\in\{0,1\}^{n}$ be uniformly distributed, and let $G\in\{0,1\}^{m\times n}$ with $m\leq n-1$. Then,
$\Pr(GY=\{0\}^{m})=\frac{|\ker(G)|}{2^{n}}.$ (54)
Proof.
Since $Y$ is uniformly distributed over $\{0,1\}^{n}$, each vector $y\in\{0,1\}^{n}$ occurs with probability $2^{-n}$. The event $GY=\{0\}^{m}$ occurs precisely when $Y\in\ker(G)$, which is a subspace of $\{0,1\}^{n}$. Thus,
$\Pr(GY=\{0\}^{m})=\Pr(Y\in\ker(G))=\frac{|\ker(G)|}{2^{n}}.$ (55)
@@
The function defined by the Circulant extractor is a two-universal family of hash functions.
@@
The function $\mathsf{Circulant}$ forms a two-universal family of hash functions if, for all distinct $x,\tilde{x}\in\{0,1\}^{n-1}$ and for $Y\in\{0,1\}^{n}$ uniformly distributed (i.e., $p_{Y}(y)=2^{-n}$ for all $y$),
$\Pr\left(\mathsf{Circulant}(x,Y)=\mathsf{Circulant}(\tilde{x},Y)\right)\leq 2^{-m}$ (56)
holds for any $m\leq n-1$.
Define $x^{\prime}=[x,0]$ and $\tilde{x}^{\prime}=[\tilde{x},0]$. Let $d^{\prime}=x^{\prime}\oplus\tilde{x}^{\prime}$, where $\oplus$ denotes bitwise addition modulo 2, and $C_{m}$ be the $m\times n$ submatrix consisting of the first $m$ rows of $\mathrm{circ}(d^{\prime})$. Using this notation, the collision probability can be rewritten as
$\Pr\left(\mathsf{Circulant}(x,Y)=\mathsf{Circulant}(\tilde{x},Y)\right)$ $=\Pr(C_{m}Y=\{0\}^{m})=p_{C_{m}Y}(\{0\}^{m}).$ (57)
Next, $x$ and $\tilde{x}$ differ in at least one element by definition, so $d^{\prime}\neq\{0\}^{n}$. Moreover, since $x^{\prime}$ and $\tilde{x}^{\prime}$ share the same last bit due to padding, we also have $d^{\prime}\neq\{1\}^{n}$. Therefore, by Lemma 3, the submatrix $C_{m}$ has rank $m$ over $\{0,1\}$ for any $m\leq n-1$.
By the rank-nullity theorem, if a matrix has rank $m$ over $\{0,1\}$, then its kernel has dimension nmn-m, and hence size $2^{n-m}$. Therefore, since $C_{m}$ has rank $m$, its kernel has size $2^{n-m}$. Applying Lemma 4, we obtain
$p_{C_{m}Y}(\{0\}^{m})=\frac{2^{n-m}}{2^{n}}=2^{-m},$ (58)
as required. ∎
@@
A.4 $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ (and $\mathsf{Circulant}$) implementation
Concretely, our implementation of the $\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}$ extractor follows [18] with $\{A_{i}\}_{i}$ chosen to be right cyclic shift matrices, which means losing one output bit compared the optimal choice for the extractor. We also require that $n$, the input length, is prime with primitive root 2 and $X=x\neq\{0\}^{n},\{1\}^{n}$. The set of matrices $A_{0},A_{1}\ldots A_{n-1}$ are defined as the $n\times n$ matrices
$A_{0}=\begin{bmatrix}1&0&0&\ldots&\ldots&0\\ 0&1&0&\ldots&\ldots&0\\ 0&0&1&\ldots&\ldots&0\\ \vdots&&\ddots&&\vdots\\ 0&0&0&\ldots&1&0\\ 0&0&0&\ldots&0&1\\ \end{bmatrix},\quad A_{1}=\begin{bmatrix}0&1&0&\ldots&\ldots&0\\ 0&0&1&\ldots&\ldots&0\\ 0&0&0&\ldots&\ldots&0\\ \vdots&&\ddots&&\vdots\\ 0&0&0&\ldots&0&1\\ 1&0&0&\ldots&0&0\\ \end{bmatrix}$$\ldots,$ (59)
and satisfy the constraint that the sum of any subset $B\subset\{A_{i}\}_{i}$ has rank at least n1n-1, as proved in [75]. In this form, the extractor can be re-written as the matrix-vector multiplication
$\mathsf{Dodis\hskip 2.84544ptet\hskip 2.84544ptal.}(x,y)=\left(\begin{bmatrix}x_{0}&x_{1}&x_{2}&\ldots&\ldots&x_{n-1}\\ x_{n-1}&x_{0}&x_{1}&\ldots&\ldots&x_{n-2}\\ x_{n-2}&x_{n-1}&x_{0}&\ldots&\ldots&x_{n-3}\\ \vdots&&\ddots&\ddots&&\vdots\\ x_{2}&x_{3}&x_{4}&\ldots&x_{0}&x_{1}\\ x_{1}&x_{2}&x_{3}&\ldots&x_{n-1}&x_{0}\\ \end{bmatrix}\begin{bmatrix}y_{0}\\ y_{1}\\ \vdots\\ \vdots\\ y_{n-1}\end{bmatrix}\right)_{{0:m-1}},$ (60)
where $x=x_{0},x_{1},\ldots x_{n-1}$, $y=y_{0},y_{1},\ldots y_{n-1}$ and the subscript 0:m10:m-1 denotes the first $m$ elements (bits) of the matrix-vector multiplication.
Although this construction is not optimal in the sense that it loses one output bit compared to the optimal construction, it allows the overall function to be re-written as a circular convolution ([18], Appendix D.1, Definition 10). This means that we can write
@@
where subscript $i\in\{0,\ldots m-1\}$ denotes the $i$th output, $\mathsf{R}(x)=x_{0},x_{n-1},x_{n-2},\ldots,x_{1}$, and indices are mod $n$. The output consists of the first $m$ elements of this convolution. The extractor algorithm is given by the pseudo-code in Algorithm 1.
@@
1:function ExtDodism,n(u, v)
2:  $L\leftarrow 2^{\text{floor}(\log_{2}(2n-2)+1)}$ $\triangleright$ Calculate NTT size